A sample of pond sediment was diluted by placing a 0.1 g into 9.9 ml of diluent, and 0.1 ml of this
dilution was spread to two Nutrient Agar (NA) plates. After 3 d incubation, 186 colonies were counted
on one plate and 198 on the other. What is the CFU/g of the sediment? (Assume that the density of the
sediment is about that of water)
Answer #1

The density of the sediment is about that of water = 1gm/ml

Therefore, when 0.1 gm of sediment was diluted into 9.9ml of diluent,

Dilution factor=0.1 gm x 1 gm/ml / 10 ml (Total volume of solution) = 0.01

CFU/gm = No of colonies / Dilution factor = 198 / 0.01 = 19800

CFU/gm = No of colonies / Dilution factor = 186 / 0.01 = 18600

If we take the average = (19800+18600) / 2 = 19200 CFU/gm